### v1

# Towards Flexible and Efficient Diffusion Low Light Enhancer

Abstract
Diffusion-based Low-Light Image Enhancement (LLIE) has demonstrated significant success in improving the visibility of low-light images. However, the substantial computational burden introduced by the iterative sampling process remains a major concern. Current acceleration methods, whether training-based or training-free, often lead to significant performance degradation. As a result, to achieve an efficient student model with performance comparable to that of existing multi-step teacher model, it is usually necessary to retrain a more capable teacher model. This approach introduces inflexibility, as it requires additional training to enhance the teacher’s performance. To address these challenges, we propose Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT), a step distillation framework specifically designed for LLIE. ReDDiT trains a student model to replicate the teacher’s trajectory in fewer steps while also possessing the ability to surpass the teacher’s performance. Specifically, we first introduce a trajectory decoder from the teacher model to provide guidance. Subsequently, a reflectance-aware trajectory refinement module is incorporated into the distillation process to enable more deterministic guidance from the teacher model. Our framework achieves comparable performance to previous diffusion-based methods with redundant steps in just 2 steps while establishing new state-of-the-art (SOTA) results with 8 or 4 steps. Comprehensive experimental evaluations on 10 benchmark datasets validate the effectiveness of our method, consistently outperforming existing SOTA methods.
Introduction
Low-light Image Enhancement (LLIE) is designed to improve the visibility and contrast of images captured in low-light conditions while maintaining natural-looking details, which is crucial for various downstream applications. Diffusion models have achieved remarkable success in this domain, demonstrating significant progress in generating photorealistic normal light images (Yi et al. 2023; Yin et al. 2023; Hou et al. 2024).
Previous diffusion-based LLIE methods primarily focus on how to condition the low-light image within the diffusion framework. Diff-Retinex applies the Retinex decomposition to the low-light image and condition these components into the diffusion model. (Yi et al. 2023). PyDiff utilizes pyramid low-light images as conditions and applies DDIM to expedite the sampling stage (Zhou, Yang, and Yang 2023; Song, Meng, and Ermon 2020). In WCDM, the diffusion model operates in the wavelet-transformed space, restoring the high-frequency parts of low-light images (Jiang et al. 2023). GSAD proposes a global structural regularization to enhance structural information learning, enabling smaller noise schedules for the inference stage (Hou et al. 2024). These methods predominantly focus on the conditioning of the LLIE and rely on universal acceleration techniques, i.e., the DDIM sampler, during the sampling,which results in poor trade-off between of the performance and efficiency.
We attempt to develop an efficient and flexible diffusion-based low-light enhancer with acceleration methods. However, a significant gap remains. Accelerating methods for diffusion models can generally be categorized into two types: training-based and training-free methods. Training-free methods, such as various ODE and SDE samplers, offer greater flexibility and convenience for compressing the sampling process. However, these methods typically maintain good performance only with 10 to 25 sampling steps. Training-based methods, including various distillation techniques (Salimans and Ho 2021; Song et al. 2023) and Rectified Flow methods (Liu 2022), can reduce the number of sampling steps to 10 through additional training. Unfortunately, these methods still suffer from significant performance degradation compared to the original model during compression, as highlighted in previous studies. This limitation implies that achieving an efficient diffusion model with comparable performance requires training a new teacher with enhanced capabilities, resulting in inflexibility.
Intuitively, if the student model can surpass the teacher during distillation, it could provide a flexible strategy for achieving efficient diffusion with strong performance. Recent distillation techniques face challenges as they operate on Gaussian noise, which introduces a larger search space. This approach leads to inefficiencies in fitting the teacher model’s trajectory, making it nearly impossible to achieve performance that exceeds the teacher model. While this issue persists in unconditional image generation, it can be addressed in LLIE since the degraded image is already available. This allows for the reduction of the search space by incorporating deterministic elements into the diffusion trajectory. The reflectance component of images, which captures the intrinsic characteristics of a scene and remains stable across varying illumination conditions, is a natural candidate for use as a deterministic prior to adjust and refine diffusion trajectories.
Motivated by this, we propose Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT). Firstly, We introduce a trajectory decoder for the teacher model to facilitate trajectory learning. Subsequently, recognizing the beneficial impact of the reflectance in low-light images, we propose a novel reflectance-aware trajectory refinement module. By incorporating illumination and noise level estimation, we predict the reflectance of low-light images based on the Retinex theory. Then the reflectance is remapped to the trajectory of the teacher diffusion model, refining its trajectory accordingly. Following this, trajectory matching distillation is conducted between the teacher and student models. To further improve the performance, auxiliary losses that directly obtain supervision signals from the data distribution are also integrated into the distillation process. Our method yields a 2-step diffusion model with comparable performance to previous 10-step diffusion-based methods, as well as 4 and 8-step diffusion models that achieve new SOTA results. In summary, our contributions are summarized as follows:
(1) We introduce ReDDiT, a novel distillation scheme to enhance the efficiency of generative diffusion models for LLIE. Notably, ReDDiT achieves high-quality image restoration in just 2 steps, setting a new standard for efficient diffusion-based models in this domain.
(2) We design a Reflectance-Aware Trajectory Refinement (RATR) module, which utilizes the reflectance component of the low-light images to refine the overall trajectory of the teacher model. This refinement process incorporate deterministic guidance to the student model, enhancing its learning effectiveness.
(3) Extensive experiments conducted on 10 benchmark datasets validate that ReDDiT consistently achieves SOTA results, showcasing its superiority in terms of both quality and efficiency, even with a minimal number of steps.
Related Work
Low-Light Image Enhancement
Enhancing low-light images is a classical task in low-level vision, with numerous solutions leveraging deep neural networks(Guo et al. 2020; Jiang et al. 2021; Wu et al. 2022). Recently, as the diffusion model has demonstrated promising results in image generation tasks, there has been growing interest in leveraging diffusion models for LLIE to achieve faithful restoration results. Diff-Retinex applies Retinex decomposition as the condition of diffusion model(Yi et al. 2023). PyDiff utilizes pyramid low-light images as the condition and applies DDIM to speed up the sampling stage (Zhou, Yang, and Yang 2023; Song, Meng, and Ermon 2020). In WCDM, the diffusion model is constructed in the wavelet-transformed space and restores the high-frequency parts of low-light images (Jiang et al. 2023). GSAD proposes a global structural regularization to enhance structural information learning, which also helps to train diffusion models with less curvature, enabling smaller noise schedules for the inference stage (Hou et al. 2024). Unfortunately, these works only consider how to condition the low-light images, neglecting the sampling strategy and raising efficiency concerns.
Diffusion Model Acceleration
Diffusion models have shown promising performance across various tasks such as style transfer (Chen et al. 2023), video generation (Deng et al. 2023), and text-to-image generation (Li et al. 2024; Ruiz et al. 2023). However, the issue of redundant sampling steps remains a significant efficiency concern. Recent research has explored various strategies to accelerate diffusion models. One series of methods involves the use of fast post-hoc samplers (Lu et al. 2022a, b; Song, Meng, and Ermon 2020), which reduces the number of inference steps for pre-trained diffusion models to 20-50 steps. However, most of them suffer from severe performance degradation when further accelerate sampling within 10 steps. To address this limitation, step distillation is proposed, aiming to distill diffusion models into fewer steps (Luhman and Luhman 2021; Salimans and Ho 2021). Progressive distillation (PD) (Salimans and Ho 2021) is the first successful practice and produces 2 step unconditional diffusion. Following PD, Meng et al. apply PD to the large-scale Stable Diffusion (Meng et al. 2023), achieving text-to-image generation with 2 steps. Consistency Distillation (CD) (Song et al. 2023) aims to learn consistency among diffusion timesteps of the teacher models (Song et al. 2023). Following this line of research, CTM extends such consistency from individual timesteps to an entire trajectory along the diffusion model, enabling faithful unconditional generation(Kim et al. 2023). Despite the booming development of diffusion distillation in other fields, techniques for LLIE are still left blank. In this paper, we present the first distilled diffusion tailored for LLIE, which will be introduced in detail in the following sections.
Figure 1: Pipiline of our proposed ReDDiT. The distillation process involves two parts: teacher model leverages the estimated reflectance to refine its trajectory and student model’s trajectory is guided by teacher’s trajectory, via a distillation loss. TD denotes the Trajectory Decoder while RATR denotes the Reflectance-Aware Trajectory Refinement.
Reflectance-Aware Diffusion Distillation
To address the challenges encountered in previous diffusion-based LLIE models, we propose ReDDiT, aiming to train a student diffusion model with parameters $\theta$ to learn the trajectory of the teacher model $\eta$ in fewer steps. In this section, we first introduce a trajectory decoder to represent the trajectory from the teacher model. Next, to incorporate the deterministic prior, we design a Reflectance-Aware Trajectory Refinement Module to refine the trajectory of the teacher model before performing distillation. In order to train a student model that surpasses the teacher model, we integrate auxiliary losses that receive direct supervision signals from the clean images during distillation. These aspects will be detailed below.
Trajectory Distillation
The essence of diffusion distillation differs from traditional knowledge distillation is that distilled diffusion models learn the trajectory from the teacher, preserving the iterative sampling characteristics unique to diffusion models. We define $G(x_{t},y,t,s)$ as the trajectory decoder that transition from time step $t$ to $s$ $(s<t)$ to represent the diffusion model trajectory:
[…]
where $u\in[s,t)$ determines intermediate steps. Draw inspiration from the denoising diffusion implicit model, we represent the decoder with the teacher neural networks $\epsilon_{\eta}$ as:
$G_{\eta}(x_{t},y,t,s)$ $=\bar{\alpha}_{s}\epsilon_{\eta}(x_{t},y,t)+\sigma_{s}\frac{x_{t}-\bar{\alpha}_{t}\epsilon_{\eta}(x_{t},y,t)}{\sigma_{t}}$ (2)
$=\frac{\sigma_{s}}{\sigma_{t}}x_{t}+(\bar{\alpha}_{s}-\frac{\sigma_{s}}{\sigma_{t}}\bar{\alpha}_{t})\epsilon_{\eta}(x_{t},y,t),$
with this, we can decode the trajectory of the teacher model from time step $s$ to $t$. For a more precise estimation of the entire trajectory, we also utilize the intermediate step $u\in[s,t)$ to estimate the trajectory, which formulates the trajectory decoder in the second order as:
$x^{\eta}_{s,u,t}$ $=G_{\eta}(G_{\eta}(x_{t},y,t,u),y,u,s),$ (3)
with a perfectly pretrained teacher with parameters $\eta$, the trajectory decoder can accurately construct the trajectory. To facilitate trajectory learning during distillation, the student model should match the trajectory decoder from $t$ to $s$. The student model then learns to match this trajectory during distillation, represented as $x^{\theta}_{s,t}=G_{\eta}(x_{t},y,t,s)$, the distillation regularization is formulated as:
$\mathcal{L}_{distill}=\|x^{\theta}_{s,t}-x^{\eta}_{s,u,t}\|_{2}^{2},$ (4)
with such regularization, the information intermediate step $u\in[s,t)$ will be distilled in student model. In practice, we utilize the clean image predicted by the trajectory both on the student and teacher to perform distillation for stable training. The predicted clean images of the teacher and student denoted as $x_{target}$ and $x_{est}$, are formulated as follows:
$x_{target}$ $=\frac{x^{\eta}_{s,u,t}-(\sigma_{s}/\sigma_{t})x_{t}}{\bar{\alpha}_{s}-(\sigma_{s}/\sigma_{t})\bar{\alpha}_{t}},\quad x_{est}=\frac{x^{\theta}_{s,t}-(\sigma_{s}/\sigma_{t})x_{t}}{\bar{\alpha}_{s}-(\sigma_{s}/\sigma_{t})\bar{\alpha}_{t}},$ (5)
and the distillation regularization is modified as:
$\mathcal{L}_{distill}=\lambda(t)\|x_{target}-x_{est}\|_{2}^{2},$ (6)
where $\lambda(t)$ is an adaptive weight that controls the distillation according to the importance schedule. Drawing inspiration from the vanilla DDPM (Ho, Jain, and Abbeel 2020), the $\lambda(t)$ is set as $\omega(t)=\max(1,\frac{\bar{\alpha}_{t}^{2}}{\sigma_{t}^{2}})$.
Figure 2: Qualitative results on LOLv1 (top) , LOLv2-real (middle), and LOLv2-synthetic (bottom). Patches highlighted in each image by the red box indicate that ReDDiT effectively enhances the visibility, preserves the color and generates finer details in normal light images. Zoom in to better observe the image details.
Reflectance-Aware Trajectory Refinement
Ideally, at every step along the trajectory, the predicted clean image should remain consistent. However, because diffusion models are designed for Gaussian flow, it is challenging for them to maintain consistent prediction results throughout the trajectory, leading to discrepancies during the student’s learning process. Fortunately, given an illumination map $h$ and the noise $z$, the reflectance that shares characteristics with the clean image can be obtained through $x=\frac{y-z}{h}$ based on the Retinex theory. The reflectance information derived from the low-light image is particularly well-suited for incorporating deterministic elements into the teacher model’s trajectory. Building on this insight, we propose the Reflectance-Aware Trajectory Refinement (RATR) module to refine the teacher’s trajectory, thereby reducing fitting errors and providing deterministic guidance to the student model
For illumination estimation, we employ the maximum channel of the low-light image $y$ to represent the estimated illumination map $h^{\prime}$, as the common practice. Regarding ISO noise estimation, similar to the previous non-learning-based denoising method (Buades, Coll, and Morel 2005), the noise can be modeled as the distance between the noisy image and the clean image. We use this distance to estimate the noise map of the input low-light images:
$z^{\prime}=|y-\psi(y)|,$ (7)
where $\psi$ represents a non-learning-based denoising operation, allowing for flexible distillation. With these estimations, we can obtain a latent clean image $\tilde{x}_{0}=\frac{y-z^{\prime}}{h^{\prime}}$. Then, we remap this latent clean image to the diffusion trajectory using the deterministic diffusion process:
$\tilde{x}_{s}=\bar{\alpha}_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}.$ (8)
Then, we refine the trajectory at timestep $s$ with a strength $w_{\text{refine}}$:
$\tilde{x}^{\eta}_{s,u,t}=x^{\eta}_{s,u,t}+w_{refine}(\tilde{x}_{s}-x^{\eta}_{s,u,t}).$ (9)
The refined trajectory is then employed to perform distillation. Extensive experiments in ablation studies show that such refinement greatly improves performance.
Auxiliary Loss
In knowledge distillation for classification, it is widely acknowledged that the student classifier often performs as well as, or even outperforms the teacher classifier. A crucial factor contributing to this success is the direct training signal derived from the data label. However, in LLIE, distillation models tend to exhibit inferior sample quality compared to the teacher. This is primarily because model optimization relies solely on the distillation loss. In ReDDiT, we extend the principles of classification distillation to our model by introducing direct signals from both pixel and feature spaces. In the pixel space, we employ L2 loss on the pixel space:
$\mathcal{L}_{pix}=\lambda_{pix}\|x_{0}-x^{\theta}_{0}(x_{t})\|_{2}^{2}.$ (10)
In the feature space, we employ the perceptual loss for enhancing the student learning in structure and texture details:
$\mathcal{L}_{per}$ $=\lambda_{per}\|\phi(x_{0})-\phi(x_{0}^{\theta}(x_{t}))\|_{2}^{2}.$ (11)
The final loss in the distillation stage integrates all the components into a single and unified training framework:
$\mathcal{L}_{total}$ $=\mathcal{L}_{distill}+\mathcal{L}_{per}+\mathcal{L}_{pix}.$ (12)
Experiments
Methods NFE LOLv1 LOLv2-real LOLv2-synthetic
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
LIME (Guo, Li, and Ling 2016) (TIP 16) 1 16.760 0.560 0.350 15.240 0.470 0.415 16.880 0.776 0.675
Zero-DCE (Guo et al. 2020) (CVPR 20) 14.861 0.562 0.335 18.059 0.580 0.313 - - -
EnlightenGAN (Jiang et al. 2021) (TIP 21) 17.483 0.652 0.322 18.640 0.677 0.309 16.570 0.734 -
RetinexNet (Wei et al. 2018) (BMVC 18) 16.770 0.462 0.474 18.371 0.723 0.365 17.130 0.798 0.754
DRBN (Yang et al. 2020) (CVPR 20) 19.860 0.834 0.155 20.130 0.830 0.147 23.220 0.927 -
KinD (Zhang, Zhang, and Guo 2019) (ACM MM 19) 20.870 0.799 0.207 17.544 0.669 0.375 16.259 0.591 0.435
KinD++ (Zhang, Zhang, and Guo 2019) (IJCV 20) 21.300 0.823 0.175 19.087 0.817 0.180 - - -
MIRNet (Zamir et al. 2022b) (TPAMI 22) 24.140 0.842 0.131 20.357 0.782 0.317 21.940 0.846 -
LLFlow (Wang et al. 2022a) (AAAI 22) 25.132 0.872 0.117 26.200 0.888 0.137 24.807 0.9193 0.067
SNR-Net (Xu et al. 2022) (CVPR 22) 26.716 0.851 0.152 27.209 0.871 0.157 27.787 0.941 0.054
LLFormer (Wang et al. 2023b) (AAAI 23) 25.758 0.823 0.167 26.197 0.819 0.209 28.006 0.927 0.061
Retinexformer (Cai et al. 2023) (ICCV 23) 27.180 0.850 - 27.710 0.856 - 29.040 0.939 -
PyDiff (Zhou, Yang, and Yang 2023) (IJCAI 23) 4 27.090 0.930 0.100 29.629 0.876 0.194 20.659 0.879 0.275
WCDM (Jiang et al. 2023) (ToG 23) 10 26.947 0.847 0.088 30.461 0.879 0.070 29.852 0.908 0.077
GSAD (Hou et al. 2024) (NIPS 23) 10 27.839 0.877 0.091 28.818 0.895 0.095 28.670 0.944 0.047
Ours-8 8 28.090 0.861 0.052 30.919 0.895 0.040 30.166 0.943 0.028
Ours-4 4 27.981 0.872 0.053 31.250 0.904 0.038 30.034 0.933 0.031
Ours-2 2 27.397 0.848 0.062 30.613 0.890 0.045 29.346 0.913 0.040
Table 1: Quantitative comparisons are conducted on the LoLv1 and LoLv2 (real and synthetic) datasets. The best result is highlighted in bold, while the second-best result is underlined. Our method significantly outperformed SOTA methods.
Figure 3: Qualitative results on NPE, VV, LIME, DICM, and MEF. Compared to Retinexformer, our method effectively enhances the visibility and preserves the color in normal light images. Zoom in to better observe the image details.
Methods SID SDSD
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
SID (Chen et al. 2018) (CVPR 18) 16.97 0.591 23.29 0.703
RetinexNet (Wei et al. 2018) (BMVC 18) 16.48 0.578 20.84 0.617
DeepUPE (Wang et al. 2019) (CVPR 19) 17.01 0.604 21.70 0.662
KinD (Zhang, Zhang, and Guo 2019) (ACM MM 19) 18.02 0.583 21.95 0.672
3DLUT (Zeng et al. 2020) (TPAMI 20) 20.11 0.592 21.66 0.655
DRBN (Yang et al. 2020) (CVPR 20) 19.02 0.577 24.08 0.868
EnGAN (Jiang et al. 2021) (TIP 21) 17.23 0.543 20.02 0.604
Uformer (Wang et al. 2022b) (CVPR 22) 18.54 0.577 23.17 0.859
Restormer (Zamir et al. 2022a) (CVPR 22) 22.27 0.649 25.67 0.827
MIRNet (Zamir et al. 2022b) (TPAMI 22) 20.84 0.605 24.38 0.864
SNR-Net (Xu et al. 2022) (CVPR 22) 22.87 0.625 29.44 0.894
Retinexformer (Cai et al. 2023) (ICCV 23) 24.44 0.680 29.77 0.896
Ours-8 25.32 0.802 29.95 0.883
Ours-4 25.85 0.792 29.52 0.857
Ours-2 25.88 0.745 28.10 0.731
Table 2: Quantitative comparisons on SID and SDSD datasets. The best result is highlighted in bold, while the second-best result is underlined.
Dataset and Implementation Details
Datasets. We evaluate the performance of our framework across various datasets showcasing noise in low-light image regions, including LOLv1 (Wei et al. 2018), LOLv2 (Yang et al. 2021), SID (Chen et al. 2018), SDSD (Wang et al. 2021), DICM (Lee, Lee, and Kim 2012), LIME (Guo, Li, and Ling 2016), MEF (Li et al. 2018), NPE (Wang et al. 2013), and VV.
LOLv1 comprises 485 pairs of low/normal-light images for training and 15 pairs for testing, taken under different exposure conditions. LOLv2 is divided into LOLv2-real and LOLv2-synthetic subsets. LOLv2-real contains 689 pairs for training and 100 for testing, adjusted for exposure time and ISO. LOLv2-Synthetic is created by analyzing the illumination distribution of low-light images, comprising 900 pairs for training and 100 for testing.
SID. The subset of the SID dataset captured by a Sony camera is adopted for evaluation. There are 2697 short-/long-exposure RAW image pairs. The low-/normal-light RGB images are obtained by using the same in-camera signal processing of SID (Chen et al. 2018) to transfer RAW to sRGB. 2099 and 598 image pairs are used for training and testing. SDSD. We adopt the static version of SDSD, captured by a Canon EOS 6D Mark II camera with an ND filter. We use 62:6 and 116:10 low-/normal-light video pairs for training and testing on indoor subsets.
DICM, LIME, MEF, NPE, and VV. We evaluate the generalization capability of our models trained on the LOLv1 and LOLv2 datasets by testing them on the DICM, LIME, MEF, NPE, and VV datasets. Since these datasets lack ground truth images, evaluation relies on visual comparison and no-reference image quality assessment methods.
Evaluation metrics. We comprehensively evaluate various Low-Light Image Enhancement (LLIE) methods, employing both full-reference and non-reference image quality metrics. In cases with paired data, we measure the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM), along with the learned perceptual image patch similarity (LPIPS). For datasets such as DICM, LIME, MEF, NPE, and VV, which lack paired data, we rely on the Naturalness Image Quality Evaluator (NIQE) for assessment.
Methods DICM LIME MEF NPE VV
LIME (Guo, Li, and Ling 2016) 5.42 3.84 3.73 4.37 3.63
Zero-DCE (Guo et al. 2020) 3.77 3.61 4.12 6.19 4.81
EnGAN (Jiang et al. 2021) 3.98 4.49 4.95 3.42 3.93
KinD++ (Zhang et al. 2021) 6.19 4.24 3.93 3.44 3.72
SCI (Ma et al. 2022) 3.73 3.52 4.07 6.44 4.05
DCC-Net (Zhang et al. 2022) 3.84 6.27 6.04 7.97 4.66
Retinexformer (Cai et al. 2023) 4.47 3.96 4.12 4.28 4.53
WCDM (Jiang et al. 2023) 6.38 4.62 4.86 4.65 5.39
GSAD (Hou et al. 2024) 4.35 3.97 4.08 5.35 3.18
Ours 3.62 3.45 3.93 3.24 3.00
Table 3: Quantitative comparisons of different methods on DICM, LIME, MEF, NPE, and VV datasets in terms of NIQE, lower values indicate better performance. The best result is highlighted in bold.
Figure 4: Quantitative comparisons on the LoLv1 and LoLv2 (Real and Synthetic) datasets with other accelerate methods. Popular acceleration methods experience a significant decline in performance as the number of steps decreases. Our method outperforms other acceleration techniques as well as the original teacher model (in terms of DDIM performance).
Comparison with State-of-the-Art Methods
Results on LOLv1 and LOLv2. On LOLv1 and LOLv2, we compare ReDDiT against LIME (Guo, Li, and Ling 2016), RetinexNet (Wei et al. 2018), KinD (Zhang, Zhang, and Guo 2019), Zero-DCE (Guo et al. 2020), DRBN (Yang et al. 2020), KinD++ (Zhang et al. 2021), EnlightenGAN (Jiang et al. 2021), MIRNet (Zamir et al. 2022b), LLFlow (Wang et al. 2022a), DCC-Net (Zhang et al. 2022), SNR-Net (Xu et al. 2022), LLFormer (Wang et al. 2023b), PyDiff (Zhou, Yang, and Yang 2023), WCDM (Wang et al. 2023a), and GSAD (Hou et al. 2024). Table 1 presents the quantitative results of various LLIE methods, showcasing our method’s superior performance across all compared metrics, including PSNR, SSIM, and LPIPS. New SOTA PSNR results of 28.090, 31.250 and 30.166 can be observed on LoLv1, LoLv2-real and LoLv2-synthetic individually. Notably, ReDDiT consistently outperforms other methods, exhibiting substantial improvements in LPIPS scores, indicative of enhanced perceptual quality. Remarkably, ReDDiT achieves SOTA performance on LOLv2-real/LOLv2-synthetic datasets across all distilled models (8, 4, 2 steps). On LOLv1 datasets, our method attains SOTA results with 8 and 4 steps sampling and remains comparable to previous methods with 2 steps. The visual comparison presented in Figure 2 further highlights the effectiveness of ReDDiT in mitigating artifacts and enhancing image details. Notably, as depicted in the red box of Figure 2, ReDDiT excels in restoring clear structures and intricate details. This underscores our method’s ability to leverage generative modeling to capture natural image distributions and retain such characteristics in the distilled models, resulting in superior visual effects.
Results on SID and SDSD. On SID and SDSD, we compare ReDDiT with SID (Chen et al. 2018), DeepUPE (Wang et al. 2019), Uformer (Wang et al. 2022b), RetinexNet (Wei et al. 2018), KinD (Zhang, Zhang, and Guo 2019), DRBN (Yang et al. 2020), EnlightenGAN (Jiang et al. 2021), MIRNet (Zamir et al. 2022b), SNR-Net (Xu et al. 2022), and Retinexformer (Cai et al. 2023). Table 2 presents the quantitative results on the SID and SDSD, indicating its capability of ReDDiT to handle more complex low-light conditions effectively. Specifically, our method establishes new SOTA PSNR values of 25.32 dB/29.95 dB on SID and SDSD dataset. The visual comparison is provided in supplementary materials.
Results on DICM, LIME, MEF, NPE, and VV. As shown in Figure 3, we directly apply the model trained on the LOLv2-synthetic dataset to these unpaired real-world datasets. From Table 3, it is evident that ReDDiT outperforms all competitors in terms of NIQE scores, demonstrating its robust generalization ability. Clearly, our model adeptly adjusts illumination conditions, effectively enhancing visibility in low-light areas while avoiding over-exposure. This successful balance confirms the superior efficacy of our proposed method.
Comparison with Other Accelerate Methods
Figure 5: The ablation study on the main components of our method. It can be observed that the refinement module significantly contributes to the final performance, with its impact becoming more pronounced as the number of sampling steps decreases.
Figure 6: The ablation study on the weights of the reflectance-aware trajectory refinement. The horizontal axis represents the strength of the refinement and the vertical axis represents PSNR values.
We conduct a comprehensive evaluation of various acceleration techniques applied to our pre-trained diffusion model. The quantitative results of this comparative analysis are presented in Figure 4. Notably, our investigation highlights the critical importance of employing distillation strategies in enhancing model performance.
Among the evaluated techniques, our distillation strategy emerges as the top performer across all datasets and all sampling steps. We achieve superior results compared to the teacher model, as demonstrated by the performance improvements over the straightforward application of DDIM. However, it is noteworthy that recent advancements in consistency distillation techniques showcase slight superiority over traditional ODE solver methods, further underscoring the significance of distillation approaches in optimizing model performance.
Most notably, ReDDiT significantly outperforms alternative techniques, demonstrating the substantial benefits of investing in additional training costs to achieve superior results. This emphasizes the pivotal role of distillation strategies in effectively transferring knowledge from a well-prepared teacher model to its student counterpart, thereby elevating the overall performance of ReDDiT. These findings underscore the significance of ReDDiT’s superior performance, reinforcing its efficacy and the value of employing sophisticated distillation techniques in advancing the SOTA in low-light image enhancement.
Ablation Study
The ablation on refinement module. The performance degradation without the refinement module is evident in Figure 5, where the absence of this module results in the most significant decline compared to the SOTA performance. Further analysis in Figure 6 delves into the impact of the refinement module in more detail.
The refinement module’s influence becomes more pronounced as the number of sampling steps decreases. In the case of 2-step distillation, Figure 6 illustrates a consistent performance improvement with increasing strength of the refinements. On LOLv2-real, the 4-step distillation results even surpass the performance of the 8-step diffusion model and achieve new SOTA results with $w_{refine}=0.8$. This trend is attributed to the heightened significance of the refinement module in addressing the increasing fitting error of the teacher model as the number of steps decreases. Therefore, the refinement module becomes crucial for accurately estimating the trajectory under such conditions.
In our previous experiments, we select the refinement strength that yielded the best performance in the previous ablation studies. This decision ensures that the refinement module effectively mitigates the fitting error of the teacher model, thereby enhancing the performance of the student model during distillation.
The ablation on the auxiliary loss. As depicted in Figure 5, conducting ablation studies on $\mathcal{L}_{pix}$ and $\mathcal{L}_{per}$ results in a slight decrease in distillation performance. This observation underscores the effectiveness of directly obtaining supervision signals from ground truth data. Interestingly, $\mathcal{L}_{per}$ exhibits a more pronounced influence on performance compared to $\mathcal{L}_{pix}$, emphasizing the significance of leveraging supervision signals in the feature space.
Furthermore, the degradation in performance resulting from ablation of either loss term is minor compared to the substantial degradation observed when the Reflectance-Aware Trajectory Refinement Module is removed. This finding highlights the critical role of the Reflectance-Aware Trajectory Refinement Module within our framework.
Efficiency Comparison
We evaluate the efficiency of our method in terms of inference time, frames per second (FPS), and number of parameters, comparing it with recent diffusion-based LLIE methods. The quantitative results of this comparative analysis are presented in Table 4. Our approach, particularly in its 2-step variant, demonstrates superior performance regarding inference speed, FPS, and parameter efficiency. It achieves an optimal balance between computational efficiency and model performance, surpassing other methods in both speed and resource utilization. This comparison highlights the effectiveness of our method in delivering high-performance results with minimal computational overhead
Teacher-16 PyDiff GSAD WCDM Ours-8 Ours-4 Ours-2
Time (s) 0.611 0.254 0.381 0.117 0.323 0.162 0.076
FPS 1.64 3.92 2.62 8.51 3.09 6.15 13.1
Param (M) 17.43 97.89 17.43 22.08 17.43 17.43 17.43
Table 4: The efficiency comparison with other methods. Our method resulted in a fast and lightweight diffusion-based low-light enhancer with excellent performance.
Limitation and Future work
While ReDDiT has excelled in restoration within 2 steps, single-step restoration is not yet optimal, presenting a notable limitation. In single-step restoration, inefficient denoising results in low PSNR and artifacts in detail. Failure cases in the single step will be demonstrated in the supplementary materials. Additionally, the exploration of a lightweight denoising network was not addressed in this work. In our future work, we will continue exploring efficient diffusion-based methods for LLIE. Our central focus will be on investigating the potential of a single-step diffusion model and developing a lightweight denoising network.
Conclusions
In this paper, we introduce ReDDiT, a significant advancement in efficient diffusion models for low-light image enhancement (LLIE). Central to ReDDiT is an innovative distillation technique that integrates deterministic elements and reduces trajectory fitting errors through Reflectance-Aware Trajectory Refinement. These advancements enable ReDDiT to preserve the intrinsic structural integrity of images while reducing sampling steps. ReDDiT achieves performance comparable to previous methods using just 2 steps and sets new state-of-the-art (SOTA) results in 4 and 8 steps. Across 10 benchmark datasets, ReDDiT consistently outperforms existing methods. This represents a promising step toward real-time diffusion models for LLIE, and our research in this area will continue.

### v2

@@
-Towards Flexible and Efficient Diffusion Low Light Enhancer
+Efficient Diffusion as Low Light Enhancer
 Abstract
-Diffusion-based Low-Light Image Enhancement (LLIE) has demonstrated significant success in improving the visibility of low-light images.
-However, the substantial computational burden introduced by the iterative sampling process remains a major concern.
-Current acceleration methods, whether training-based or training-free, often lead to significant performance degradation.
-As a result, to achieve an efficient student model with performance comparable to that of existing multi-step teacher model, it is usually necessary to retrain a more capable teacher model.
-This approach introduces inflexibility, as it requires additional training to enhance the teacher’s performance.
-To address these challenges, we propose Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT), a step distillation framework specifically designed for LLIE.
-ReDDiT trains a student model to replicate the teacher’s trajectory in fewer steps while also possessing the ability to surpass the teacher’s performance.
-Specifically, we first introduce a trajectory decoder from the teacher model to provide guidance.
-Subsequently, a reflectance-aware trajectory refinement module is incorporated into the distillation process to enable more deterministic guidance from the teacher model.
+The computational burden of the iterative sampling process remains a major challenge in diffusion-based Low-Light Image Enhancement (LLIE).
+Current acceleration methods, whether training-based or training-free, often lead to significant performance degradation, highlighting the trade-off between performance and efficiency.
+In this paper, we identify two primary factors contributing to performance degradation: fitting errors and the inference gap.
+Our key insight is that fitting errors can be mitigated by linearly extrapolating the incorrect score functions, while the inference gap can be reduced by shifting the Gaussian flow to a reflectance-aware residual space.
+Based on the above insights, we design Reflectance-Aware Trajectory Refinement (RATR) module, a simple yet effective module to refine the teacher trajectory using the reflectance component of images.
+Following this, we introduce Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT), an efficient and flexible distillation framework tailored for LLIE.
 Our framework achieves comparable performance to previous diffusion-based methods with redundant steps in just 2 steps while establishing new state-of-the-art (SOTA) results with 8 or 4 steps.
 Comprehensive experimental evaluations on 10 benchmark datasets validate the effectiveness of our method, consistently outperforming existing SOTA methods.
-Introduction
+[…]
+1 Introduction
+Figure 1: ReDDiT shifts the teacher trajectory from the original Gaussian distribution to a residual space, effectively reducing the sampling gap.
+Subsequently, it refines the teacher trajectory towards the ground truth trajectory to mitigate fitting errors.
 Low-light Image Enhancement (LLIE) is designed to improve the visibility and contrast of images captured in low-light conditions while maintaining natural-looking details, which is crucial for various downstream applications.
-Diffusion models have achieved remarkable success in this domain, demonstrating significant progress in generating photorealistic normal light images (Yi et al.
-2023; Yin et al.
-2023; Hou et al.
-2024).
-Previous diffusion-based LLIE methods primarily focus on how to condition the low-light image within the diffusion framework.
-Diff-Retinex applies the Retinex decomposition to the low-light image and condition these components into the diffusion model.
-(Yi et al.
-2023).
-PyDiff utilizes pyramid low-light images as conditions and applies DDIM to expedite the sampling stage (Zhou, Yang, and Yang 2023; Song, Meng, and Ermon 2020).
-In WCDM, the diffusion model operates in the wavelet-transformed space, restoring the high-frequency parts of low-light images (Jiang et al.
-2023).
-GSAD proposes a global structural regularization to enhance structural information learning, enabling smaller noise schedules for the inference stage (Hou et al.
-2024).
-These methods predominantly focus on the conditioning of the LLIE and rely on universal acceleration techniques, i.e., the DDIM sampler, during the sampling,which results in poor trade-off between of the performance and efficiency.
-We attempt to develop an efficient and flexible diffusion-based low-light enhancer with acceleration methods.
-However, a significant gap remains.
-Accelerating methods for diffusion models can generally be categorized into two types: training-based and training-free methods.
-Training-free methods, such as various ODE and SDE samplers, offer greater flexibility and convenience for compressing the sampling process.
-However, these methods typically maintain good performance only with 10 to 25 sampling steps.
-Training-based methods, including various distillation techniques (Salimans and Ho 2021; Song et al.
-2023) and Rectified Flow methods (Liu 2022), can reduce the number of sampling steps to 10 through additional training.
-Unfortunately, these methods still suffer from significant performance degradation compared to the original model during compression, as highlighted in previous studies.
-This limitation implies that achieving an efficient diffusion model with comparable performance requires training a new teacher with enhanced capabilities, resulting in inflexibility.
-Intuitively, if the student model can surpass the teacher during distillation, it could provide a flexible strategy for achieving efficient diffusion with strong performance.
-Recent distillation techniques face challenges as they operate on Gaussian noise, which introduces a larger search space.
-This approach leads to inefficiencies in fitting the teacher model’s trajectory, making it nearly impossible to achieve performance that exceeds the teacher model.
-While this issue persists in unconditional image generation, it can be addressed in LLIE since the degraded image is already available.
-This allows for the reduction of the search space by incorporating deterministic elements into the diffusion trajectory.
-The reflectance component of images, which captures the intrinsic characteristics of a scene and remains stable across varying illumination conditions, is a natural candidate for use as a deterministic prior to adjust and refine diffusion trajectories.
-Motivated by this, we propose Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT).
-Firstly, We introduce a trajectory decoder for the teacher model to facilitate trajectory learning.
-Subsequently, recognizing the beneficial impact of the reflectance in low-light images, we propose a novel reflectance-aware trajectory refinement module.
-By incorporating illumination and noise level estimation, we predict the reflectance of low-light images based on the Retinex theory.
-Then the reflectance is remapped to the trajectory of the teacher diffusion model, refining its trajectory accordingly.
-Following this, trajectory matching distillation is conducted between the teacher and student models.
-To further improve the performance, auxiliary losses that directly obtain supervision signals from the data distribution are also integrated into the distillation process.
-Our method yields a 2-step diffusion model with comparable performance to previous 10-step diffusion-based methods, as well as 4 and 8-step diffusion models that achieve new SOTA results.
-In summary, our contributions are summarized as follows:
-(1) We introduce ReDDiT, a novel distillation scheme to enhance the efficiency of generative diffusion models for LLIE.
-Notably, ReDDiT achieves high-quality image restoration in just 2 steps, setting a new standard for efficient diffusion-based models in this domain.
-(2) We design a Reflectance-Aware Trajectory Refinement (RATR) module, which utilizes the reflectance component of the low-light images to refine the overall trajectory of the teacher model.
-This refinement process incorporate deterministic guidance to the student model, enhancing its learning effectiveness.
+Diffusion models [9, 22] have achieved remarkable success in this domain, demonstrating significant progress in generating photorealistic normal light images [40, 41, 10].
+Previous diffusion-based LLIE methods [40, 41, 10] primarily focus on how to condition the low-light image within the diffusion framework.
+Diff-Retinex [40] applies the Retinex decomposition to the low-light image and condition these components into the diffusion model.
+PyDiff [48] utilizes pyramid low-light images as conditions and applies DDIM [26] to expedite the sampling stage.
+In WCDM [11], the diffusion model operates in the wavelet-transformed space, restoring the high-frequency parts of low-light images.
+GSAD [10] proposes a global structural regularization to enhance structural information learning, enabling smaller noise schedules for the inference stage.
+However, the primary challenge to the broader application of diffusion models in LLIE stems from their iterative denoising mechanism.
+For example, DDPM [9] requires multiple denoising steps, e.g., 1000 steps, to reverse a Gaussian noise into a clean image.
+The computation overhead conflicts with the demands of LLIE, particularly in computational photography applications for edge devices like mobile phones and surveillance cameras.
+Inspired by training-aware acceleration methods for diffusion models [25, 27], we investigate distilling a multi-step diffusion-based LLIE model to improve efficiency.
+A key observation is that performance degradation is inevitable as the number of sampling steps is reduced, even when using advanced acceleration methods (e.g., consistency distillation [27]).
+This leads us to pose the question: Is it possible to distill a student diffusion model that surpasses the original diffusion model?
+If so, even as performance degrades with fewer sampling steps, we could still achieve results comparable to those of the multi-step teacher model.
+To address this, we conduct a comprehensive analysis of diffusion model acceleration techniques.
+Through theoretical analysis, we identify two primary factors contributing to performance degradation: (1) The fitting error.
+It is the unavoidable errors between deep learning models and target fitting data.
+It will result the additional undesired terms and mismatch during distillation.
+(2) The inference gap.
+It is the gap between the training target and sampling strategies specialized for the Diffusion models.
+It is caused by the universal diffusion model often trained and operated on the Gaussian flow for the diversity of generation, while the LLIE requires for more deterministic.
+Our key insight is to refine the teacher trajectory by applying linear extrapolation to the score functions of the teacher model, thereby mitigating the adverse effects of fitting errors.
+Meanwhile, shifting the sampling trajectory to a deterministic space addresses the sampling gap, as shown in Figure 1.
+A detailed theoretical analysis of this approach is presented in  Sec.
+3.
+Based on the above principles, we design Reflectance-Aware Trajectory Refinement (RATR) module, a simple yet effective module to refine the teacher trajectory for LLIE task.
+It incorporates the reflectance component of images as a deterministic prior to adjust and refine diffusion trajectories.
+Following this, we introduce Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT), an efficient and flexible distillation framework tailored for LLIE.
+This framework perform trajectory matching distillation is conducted between the refined teacher and student models, yields a 2-step diffusion model with comparable performance to previous 10-step diffusion-based methods, as well as 4 and 8-step diffusion models that achieve new SOTA results.
+our contributions are summarized as follows:
+(1) We theoretically analyze the factors contributing to performance degradation and propose targeted improvements: linear extrapolation of the score function to mitigate performance loss due to fitting errors, and residual shifting as a solution to the sampling gap.
+(2) Based on these two design principles, we present ReDDiT, a novel distillation scheme that enhances the efficiency of generative diffusion models for LLIE.
+Notably, ReDDiT achieves high-quality image restoration in just 2 steps, establishing a new benchmark for efficient diffusion-based models in this domain.
 (3) Extensive experiments conducted on 10 benchmark datasets validate that ReDDiT consistently achieves SOTA results, showcasing its superiority in terms of both quality and efficiency, even with a minimal number of steps.
-Related Work
-Low-Light Image Enhancement
-Enhancing low-light images is a classical task in low-level vision, with numerous solutions leveraging deep neural networks(Guo et al.
-2020; Jiang et al.
-2021; Wu et al.
-2022).
+2 Related Work
+Low-Light Image Enhancement.
+Enhancing low-light images is a classical task in low-level vision, with numerous solutions leveraging deep neural networks [7, 12, 36].
 Recently, as the diffusion model has demonstrated promising results in image generation tasks, there has been growing interest in leveraging diffusion models for LLIE to achieve faithful restoration results.
-Diff-Retinex applies Retinex decomposition as the condition of diffusion model(Yi et al.
-2023).
-PyDiff utilizes pyramid low-light images as the condition and applies DDIM to speed up the sampling stage (Zhou, Yang, and Yang 2023; Song, Meng, and Ermon 2020).
-In WCDM, the diffusion model is constructed in the wavelet-transformed space and restores the high-frequency parts of low-light images (Jiang et al.
-2023).
-GSAD proposes a global structural regularization to enhance structural information learning, which also helps to train diffusion models with less curvature, enabling smaller noise schedules for the inference stage (Hou et al.
-2024).
-Unfortunately, these works only consider how to condition the low-light images, neglecting the sampling strategy and raising efficiency concerns.
-Diffusion Model Acceleration
-Diffusion models have shown promising performance across various tasks such as style transfer (Chen et al.
-2023), video generation (Deng et al.
-2023), and text-to-image generation (Li et al.
-2024; Ruiz et al.
-2023).
+Diff-Retinex applies Retinex decomposition as the condition of diffusion model [40].
+PyDiff utilizes pyramid low-light images as the condition and applies DDIM to speed up the sampling stage [48, 26].
+In WCDM, the diffusion model is constructed in the wavelet-transformed space and restores the high-frequency parts of low-light images [11].
+GSAD proposes a global structural regularization to enhance structural information learning, which also helps to train diffusion models with less curvature, enabling smaller noise schedules for the inference stage [10].
+Unfortunately, these works only consider how to condition the low-light images, neglecting the efficiency concerns.
+Diffusion Model Acceleration.
+Diffusion models have shown promising performance across various tasks such as style transfer [4], video generation [5], and text-to-image generation [16, 23, 22].
 However, the issue of redundant sampling steps remains a significant efficiency concern.
 Recent research has explored various strategies to accelerate diffusion models.
-One series of methods involves the use of fast post-hoc samplers (Lu et al.
-2022a, b; Song, Meng, and Ermon 2020), which reduces the number of inference steps for pre-trained diffusion models to 20-50 steps.
-However, most of them suffer from severe performance degradation when further accelerate sampling within 10 steps.
-To address this limitation, step distillation is proposed, aiming to distill diffusion models into fewer steps (Luhman and Luhman 2021; Salimans and Ho 2021).
-Progressive distillation (PD) (Salimans and Ho 2021) is the first successful practice and produces 2 step unconditional diffusion.
-Following PD, Meng et al.
-apply PD to the large-scale Stable Diffusion (Meng et al.
-2023), achieving text-to-image generation with 2 steps.
-Consistency Distillation (CD) (Song et al.
-2023) aims to learn consistency among diffusion timesteps of the teacher models (Song et al.
-2023).
-Following this line of research, CTM extends such consistency from individual timesteps to an entire trajectory along the diffusion model, enabling faithful unconditional generation(Kim et al.
-2023).
+One series of methods involves the use of fast post-hoc samplers [17, 18, 26], which reduces the number of inference steps for pre-trained diffusion models to 20-50 steps.
+However, most suffer from severe performance degradation when further accelerating sampling within 10 steps.
+To address this limitation, step distillation [19, 25] is proposed, aiming to distill diffusion models into fewer steps.
+Progressive distillation (PD) [25] is the first successful practice and produces 2-step unconditional diffusion.
+Following PD [25], Meng et al.
+apply PD to the large-scale Stable Diffusion, achieving text-to-image generation with 2 steps.
+Consistency Distillation (CD) [27] aims to learn consistency among diffusion timesteps of the teacher models [27].
+Following this line of research, CTM [13] extends such consistency from individual timesteps to an entire trajectory along the diffusion model, enabling faithful unconditional generation.
 Despite the booming development of diffusion distillation in other fields, techniques for LLIE are still left blank.
@@
 TD denotes the Trajectory Decoder while RATR denotes the Reflectance-Aware Trajectory Refinement.
-Reflectance-Aware Diffusion Distillation
-To address the challenges encountered in previous diffusion-based LLIE models, we propose ReDDiT, aiming to train a student diffusion model with parameters $\theta$ to learn the trajectory of the teacher model $\eta$ in fewer steps.
-In this section, we first introduce a trajectory decoder to represent the trajectory from the teacher model.
-Next, to incorporate the deterministic prior, we design a Reflectance-Aware Trajectory Refinement Module to refine the trajectory of the teacher model before performing distillation.
-In order to train a student model that surpasses the teacher model, we integrate auxiliary losses that receive direct supervision signals from the clean images during distillation.
-These aspects will be detailed below.
-Trajectory Distillation
+3 Methods
+Let $x_{0}$ denote a high-quality image, the forward diffusion process aims to generate a sequence of noisy latent variables $x_{1},x_{2},...,x_{T}$ using a Markovian process, which is defined by the equation:
+$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon,$ (1)
+where $\alpha_{t}\in(0,1)$ represents the noise schedule, $\sigma_{t}$ denotes the covariance at $t$, and $\epsilon$ represents the Gaussian noise.
+In the reverse process, we incorporate the low-light image $y$ as the conditioning of the score functions $\epsilon_{\eta}(x_{t},y,\alpha_{t})$ and predict the clean images iteratively:
+$x^{\eta}_{t-1}=\alpha_{t-1}\left(\frac{x_{t}-\sigma_{t}\epsilon_{\eta}}{\alpha_{t}}\right)+\sigma_{t-1}\epsilon_{\eta}.$ (2)
+In this section, we propose ReDDiT, aiming to train a student diffusion model with parameters $\theta$ to learn the trajectory of the teacher model $\eta$ in fewer steps.
+We begin by introducing trajectory distillation.
+Next, we theoretically analyze the factors contributing to performance degradation and present a concise formulation for refining the trajectory.
+Then, we determine that the reflectance component is most effective for refining the teacher trajectory and design a Reflectance-Aware Trajectory Refinement (RATR) module.
+Finally, all components are integrated into the ReDDiT framework, where distillation is performed.
+3.1 Trajectory Distillation
 The essence of diffusion distillation differs from traditional knowledge distillation is that distilled diffusion models learn the trajectory from the teacher, preserving the iterative sampling characteristics unique to diffusion models.
-We define $G(x_{t},y,t,s)$ as the trajectory decoder that transition from time step $t$ to $s$ $(s<t)$ to represent the diffusion model trajectory:
-[…]
-where $u\in[s,t)$ determines intermediate steps. Draw inspiration from the denoising diffusion implicit model, we represent the decoder with the teacher neural networks $\epsilon_{\eta}$ as:
-$G_{\eta}(x_{t},y,t,s)$ $=\bar{\alpha}_{s}\epsilon_{\eta}(x_{t},y,t)+\sigma_{s}\frac{x_{t}-\bar{\alpha}_{t}\epsilon_{\eta}(x_{t},y,t)}{\sigma_{t}}$ (2)
-$=\frac{\sigma_{s}}{\sigma_{t}}x_{t}+(\bar{\alpha}_{s}-\frac{\sigma_{s}}{\sigma_{t}}\bar{\alpha}_{t})\epsilon_{\eta}(x_{t},y,t),$
-with this, we can decode the trajectory of the teacher model from time step $s$ to $t$.
-For a more precise estimation of the entire trajectory, we also utilize the intermediate step $u\in[s,t)$ to estimate the trajectory, which formulates the trajectory decoder in the second order as:
-$x^{\eta}_{s,u,t}$ $=G_{\eta}(G_{\eta}(x_{t},y,t,u),y,u,s),$ (3)
-with a perfectly pretrained teacher with parameters $\eta$, the trajectory decoder can accurately construct the trajectory.
-To facilitate trajectory learning during distillation, the student model should match the trajectory decoder from $t$ to $s$.
-The student model then learns to match this trajectory during distillation, represented as $x^{\theta}_{s,t}=G_{\eta}(x_{t},y,t,s)$, the distillation regularization is formulated as:
-$\mathcal{L}_{distill}=\|x^{\theta}_{s,t}-x^{\eta}_{s,u,t}\|_{2}^{2},$ (4)
+We define $G(x_{t},y,t,s)$ as the trajectory decoder that transition from time step $t$ to $s$ $(s<t)$ to represent the diffusion model trajectory.
+Draw inspiration from the denoising diffusion implicit model, we represent the decoder with the student score function $\epsilon_{\theta}$ as:
+$G_{\theta}(x_{t},y,t,s)=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\epsilon_{\theta}(x_{t},y,t).$ (3)
+With this trajectory decoder, the teacher model’s trajectory can be decoded from time step $s$ to $t$.
+For a more precise estimation of the entire teacher trajectory, we also utilize the intermediate step $u\in[s,t)$ to estimate the trajectory, which formulates the trajectory decoder in the second order as:
+$x^{\eta}_{s,u,t}$ $=G_{\eta}(G_{\eta}(x_{t},y,t,u),y,u,s)$ (4)
+$=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$
+$+\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}(\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u))$
+$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\epsilon_{\eta}(x_{t},y,t),$
+To facilitate trajectory learning during distillation, the student model should match the teacher trajectory from $t$ to $s$.
+Denoting the student trajectory as $x^{\theta}_{s,t}=G_{\eta}(x_{t},y,t,s)$, the distillation regularization is formulated as:
+$\mathcal{L}_{distill}=\|x^{\theta}_{s,t}-x^{\eta}_{s,u,t}\|_{2}^{2}.$ (5)
 With such regularization, the information intermediate step $u\in[s,t)$ will be distilled in student model.
 In practice, we utilize the clean image predicted by the trajectory both on the student and teacher to perform distillation for stable training.
-The predicted clean images of the teacher and student denoted as $x_{target}$ and $x_{est}$, are formulated as follows:
+The predicted clean images of the teacher and student, denoted as $x_{target}$ and $x_{est}$, are formulated as follows:
 $x_{target}$ $=\frac{x^{\eta}_{s,u,t}-(\sigma_{s}/\sigma_{t})x_{t}}{\bar{\alpha}_{s}-(\sigma_{s}/\sigma_{t})\bar{\alpha}_{t}},\quad x_{est}=\frac{x^{\theta}_{s,t}-(\sigma_{s}/\sigma_{t})x_{t}}{\bar{\alpha}_{s}-(\sigma_{s}/\sigma_{t})\bar{\alpha}_{t}},$ (6)
@@
 $\mathcal{L}_{distill}=\lambda(t)\|x_{target}-x_{est}\|_{2}^{2},$ (7)
-where $\lambda(t)$ is an adaptive weight that controls the distillation according to the importance schedule.
-Drawing inspiration from the vanilla DDPM (Ho, Jain, and Abbeel 2020), the $\lambda(t)$ is set as $\omega(t)=\max(1,\frac{\bar{\alpha}_{t}^{2}}{\sigma_{t}^{2}})$.
-Figure 2: Qualitative results on LOLv1 (top) , LOLv2-real (middle), and LOLv2-synthetic (bottom).
-Patches highlighted in each image by the red box indicate that ReDDiT effectively enhances the visibility, preserves the color and generates finer details in normal light images.
-Zoom in to better observe the image details.
-Reflectance-Aware Trajectory Refinement
-Ideally, at every step along the trajectory, the predicted clean image should remain consistent.
-However, because diffusion models are designed for Gaussian flow, it is challenging for them to maintain consistent prediction results throughout the trajectory, leading to discrepancies during the student’s learning process.
-Fortunately, given an illumination map $h$ and the noise $z$, the reflectance that shares characteristics with the clean image can be obtained through $x=\frac{y-z}{h}$ based on the Retinex theory.
-The reflectance information derived from the low-light image is particularly well-suited for incorporating deterministic elements into the teacher model’s trajectory.
-Building on this insight, we propose the Reflectance-Aware Trajectory Refinement (RATR) module to refine the teacher’s trajectory, thereby reducing fitting errors and providing deterministic guidance to the student model
+where $\lambda(t)$ is an adaptive weight and is set as $\max(1,\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}})$.
+3.2 On the Refinement of the Teacher Trajectory
+Direct application of trajectory distillation often leads to significant performance degradation.
+In this section, we theoretically analyze the core reasons behind this degradation and propose a strategy to mitigate its effects.
+On the fitting error of the teacher trajectory.
+To understand this, let us revisit Equation 3 and Equation 4, and consider the ideal condition where $\mathcal{L}_{distill}=0$, meaning that the student’s trajectory perfectly matches the teacher’s trajectory.
+Under this condition, we have:
+$(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\epsilon_{\theta}(x_{t},y,t)$ $=\sigma_{s}\epsilon_{\eta}(x_{u},y,u)-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\epsilon_{\eta}(x_{t},y,t)$ (8)
+$+\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}.$
+Since the teacher model is trained using the vanilla diffusion loss function, the ideal teacher trajectory satisfies the condition $\epsilon_{\eta}(x_{t},y,t)=\epsilon_{\eta}(x_{u},y,u)=\tilde{\epsilon}$.
+Under this condition, the Equation 8 can be simplified as:
+$(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\epsilon_{\theta}(x_{t},y,t)$ $=(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\tilde{\epsilon}.$ (9)
+However, the existence of fitting errors makes it impossible to achieve this goal. The presence of guidance with undesired components inevitably leads to suboptimal results.
+On the mitigation of the fitting error.
+Fortunately, we find that these undesired terms can be mitigated by a scaling parameter $\omega\in(0,1]$.
+For the term $\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}$, we can further reduce its impact by multiplying it by the scaling parameter $\omega$.
+The mismatch between $\epsilon_{\eta}(x_{u})$ and $\epsilon_{\eta}(x_{t})$ introduces a more significant optimization conflict.
+Due to the presence of fitting error, this term is not always consistent with $\epsilon_{\eta}(x_{t},y,t)$.
+Our key insight is to refine this term through linear interpolation toward the ideal value, expressed as $\sigma_{t}\Big{(}\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon}\Big{)}$.
+This approach not only maintains consistency with previous terms but also provides linear trajectory guidance.
+After applying these operations, the refined teacher trajectory can be expressed as follows:
+$\tilde{x}^{\eta}_{s,u,t}$ $=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$ (10)
+$+\omega\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}$
+$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\Big{(}\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon}\Big{)}.$
+To this end, the distillation loss, $\mathcal{L}_{distill}$, is equivalent to the following, disregarding the coefficients of each term:
+$\mathcal{L}_{distill}$ $=||\epsilon_{\theta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)||$ (11)
+$+||\epsilon_{\theta}(x_{t},y,t)-(\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon})||.$
+On the mitigation of the inference gap.
+Considering the vanilla diffusion are trained on the Gaussian flow, we determine the $\tilde{\epsilon}$ not the pure Gaussian noise but shift it into a residual space as:
+$\tilde{\epsilon}=\frac{x_{t}-\alpha_{t}\tilde{x}_{0}}{\sigma_{t}}=\frac{\alpha_{t}(x_{0}-\tilde{x}_{0})}{\sigma_{t}}+\epsilon,$ (12)
+where $\tilde{{x}}_{0}$ should lie between the low-light and clean image spaces, serving as an intermediate initial space for the student model to learn.
+This positioning ensures a closer initial distribution for the student model compared to the Gaussian distribution
+Additionally, our investigation shows that the refinement of the teacher trajectory in Equation 10 can be implemented through a straightforward approach, as outlined in the Corollary 1.
+Corollary 1 (Proof in the supplementary material).
+Given the refinement component $\tilde{x}_{s}=\alpha_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}$, the Equation 10 is equivalent to : $\tilde{x}^{\eta}_{s,u,t}=\omega x^{\eta}_{s,u,t}+(1-\omega)\tilde{x}_{s}$.
+3.3 Reflectance-Aware Trajectory Refinement
+On the determination of the component $\tilde{x}_{0}$, using the ground truth clean images $x_{0}$ will cause $\tilde{\epsilon}$ degrade to $\epsilon$.
+A natural approach is to use the low-light images $y$ as $\tilde{x}_{0}$.
+In practice, the reflectance, which shares characteristics with both the clean images and the low-light images $y$, serves as a better component for refining the trajectory.
+Building on this observation, we propose the RATR module to refine the teacher’s trajectory, thereby reducing inference gap.
+Given an illumination map $h$ and the noise $z$, the reflectance can be obtained through $x=\frac{y-z}{h}$ based on the Retinex theory.
 For illumination estimation, we employ the maximum channel of the low-light image $y$ to represent the estimated illumination map $h^{\prime}$, as the common practice.
-Regarding ISO noise estimation, similar to the previous non-learning-based denoising method (Buades, Coll, and Morel 2005), the noise can be modeled as the distance between the noisy image and the clean image.
+Regarding ISO noise estimation, similar to the previous non-learning-based denoising method [1], the noise can be modeled as the distance between the noisy image and the clean image.
 We use this distance to estimate the noise map of the input low-light images:
@@
 With these estimations, we can obtain a latent clean image $\tilde{x}_{0}=\frac{y-z^{\prime}}{h^{\prime}}$.
-Then, we remap this latent clean image to the diffusion trajectory using the deterministic diffusion process:
-$\tilde{x}_{s}=\bar{\alpha}_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}.$ (8)
-Then, we refine the trajectory at timestep $s$ with a strength $w_{\text{refine}}$:
-$\tilde{x}^{\eta}_{s,u,t}=x^{\eta}_{s,u,t}+w_{refine}(\tilde{x}_{s}-x^{\eta}_{s,u,t}).$ (9)
-The refined trajectory is then employed to perform distillation. Extensive experiments in ablation studies show that such refinement greatly improves performance.
-Auxiliary Loss
-In knowledge distillation for classification, it is widely acknowledged that the student classifier often performs as well as, or even outperforms the teacher classifier.
-A crucial factor contributing to this success is the direct training signal derived from the data label.
-However, in LLIE, distillation models tend to exhibit inferior sample quality compared to the teacher.
-This is primarily because model optimization relies solely on the distillation loss.
-In ReDDiT, we extend the principles of classification distillation to our model by introducing direct signals from both pixel and feature spaces.
+The $\tilde{x}_{0}$ is then employed to the trajectory refinement to perform distillation.
+Extensive experiments in ablation studies show that such refinement reduce the inference gap and greatly improves performance.
+Figure 3: Qualitative results on LOLv1 (top) , LOLv2-real (middle), and LOLv2-synthetic (bottom).
+Patches highlighted in each image by the red box indicate that ReDDiT effectively enhances the visibility, preserves the color and generates finer details in normal light images.
+Zoom in to better observe the image details.
+3.4 Auxiliary Loss
+In knowledge distillation for classification, direct training signal derived from the data label will help student classifier outperforms the teacher classifier.
+In ReDDiT, we extend the principles to our model by introducing direct signals from both pixel and feature spaces.
 In the pixel space, we employ L2 loss on the pixel space:
-$\mathcal{L}_{pix}=\lambda_{pix}\|x_{0}-x^{\theta}_{0}(x_{t})\|_{2}^{2}.$ (10)
+$\mathcal{L}_{pix}=\lambda_{pix}\|x_{0}-x_{est}\|_{2}^{2}.$ (14)
 In the feature space, we employ the perceptual loss for enhancing the student learning in structure and texture details:
-$\mathcal{L}_{per}$ $=\lambda_{per}\|\phi(x_{0})-\phi(x_{0}^{\theta}(x_{t}))\|_{2}^{2}.$ (11)
+$\mathcal{L}_{per}$ $=\lambda_{per}\|\phi(x_{0})-\phi(x_{est})\|_{2}^{2}.$ (15)
 The final loss in the distillation stage integrates all the components into a single and unified training framework:
 $\mathcal{L}_{total}$ $=\mathcal{L}_{distill}+\mathcal{L}_{per}+\mathcal{L}_{pix}.$ (16)
-Experiments
+4 Experiments
 Methods NFE LOLv1 LOLv2-real LOLv2-synthetic
 PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-LIME (Guo, Li, and Ling 2016) (TIP 16) 1 16.760 0.560 0.350 15.240 0.470 0.415 16.880 0.776 0.675
-Zero-DCE (Guo et al. 2020) (CVPR 20) 14.861 0.562 0.335 18.059 0.580 0.313 - - -
-EnlightenGAN (Jiang et al. 2021) (TIP 21) 17.483 0.652 0.322 18.640 0.677 0.309 16.570 0.734 -
-RetinexNet (Wei et al. 2018) (BMVC 18) 16.770 0.462 0.474 18.371 0.723 0.365 17.130 0.798 0.754
-DRBN (Yang et al. 2020) (CVPR 20) 19.860 0.834 0.155 20.130 0.830 0.147 23.220 0.927 -
-KinD (Zhang, Zhang, and Guo 2019) (ACM MM 19) 20.870 0.799 0.207 17.544 0.669 0.375 16.259 0.591 0.435
-KinD++ (Zhang, Zhang, and Guo 2019) (IJCV 20) 21.300 0.823 0.175 19.087 0.817 0.180 - - -
-MIRNet (Zamir et al. 2022b) (TPAMI 22) 24.140 0.842 0.131 20.357 0.782 0.317 21.940 0.846 -
-LLFlow (Wang et al. 2022a) (AAAI 22) 25.132 0.872 0.117 26.200 0.888 0.137 24.807 0.9193 0.067
-SNR-Net (Xu et al. 2022) (CVPR 22) 26.716 0.851 0.152 27.209 0.871 0.157 27.787 0.941 0.054
-LLFormer (Wang et al. 2023b) (AAAI 23) 25.758 0.823 0.167 26.197 0.819 0.209 28.006 0.927 0.061
-Retinexformer (Cai et al. 2023) (ICCV 23) 27.180 0.850 - 27.710 0.856 - 29.040 0.939 -
-PyDiff (Zhou, Yang, and Yang 2023) (IJCAI 23) 4 27.090 0.930 0.100 29.629 0.876 0.194 20.659 0.879 0.275
-WCDM (Jiang et al. 2023) (ToG 23) 10 26.947 0.847 0.088 30.461 0.879 0.070 29.852 0.908 0.077
-GSAD (Hou et al. 2024) (NIPS 23) 10 27.839 0.877 0.091 28.818 0.895 0.095 28.670 0.944 0.047
+LIME [8] (TIP 16) 1 16.760 0.560 0.350 15.240 0.470 0.415 16.880 0.776 0.675
+Zero-DCE [7] (CVPR 20) 14.861 0.562 0.335 18.059 0.580 0.313 - - -
+EnlightenGAN [12] (TIP 21) 17.483 0.652 0.322 18.640 0.677 0.309 16.570 0.734 -
+RetinexNet [35] (BMVC 18) 16.770 0.462 0.474 18.371 0.723 0.365 17.130 0.798 0.754
+DRBN [38] (CVPR 20) 19.860 0.834 0.155 20.130 0.830 0.147 23.220 0.927 -
+KinD [45] (ACM MM 19) 20.870 0.799 0.207 17.544 0.669 0.375 16.259 0.591 0.435
+KinD++ [45] (IJCV 20) 21.300 0.823 0.175 19.087 0.817 0.180 - - -
+MIRNet [43] (TPAMI 22) 24.140 0.842 0.131 20.357 0.782 0.317 21.940 0.846 -
+LLFlow [33] (AAAI 22) 25.132 0.872 0.117 26.200 0.888 0.137 24.807 0.9193 0.067
+SNR-Net [37] (CVPR 22) 26.716 0.851 0.152 27.209 0.871 0.157 27.787 0.941 0.054
+LLFormer [32] (AAAI 23) 25.758 0.823 0.167 26.197 0.819 0.209 28.006 0.927 0.061
+Retinexformer [2] (ICCV 23) 27.180 0.850 - 27.710 0.856 - 29.040 0.939 -
+PyDiff [48] (IJCAI 23) 4 27.090 0.930 0.100 29.629 0.876 0.194 20.659 0.879 0.275
+WCDM [11] (ToG 23) 10 26.947 0.847 0.088 30.461 0.879 0.070 29.852 0.908 0.077
+GSAD [10] (NIPS 23) 10 27.839 0.877 0.091 28.818 0.895 0.095 28.670 0.944 0.047
 Ours-8 8 28.090 0.861 0.052 30.919 0.895 0.040 30.166 0.943 0.028
@@
 PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-SID (Chen et al. 2018) (CVPR 18) 16.97 0.591 23.29 0.703
-RetinexNet (Wei et al. 2018) (BMVC 18) 16.48 0.578 20.84 0.617
-DeepUPE (Wang et al. 2019) (CVPR 19) 17.01 0.604 21.70 0.662
-KinD (Zhang, Zhang, and Guo 2019) (ACM MM 19) 18.02 0.583 21.95 0.672
-3DLUT (Zeng et al. 2020) (TPAMI 20) 20.11 0.592 21.66 0.655
-DRBN (Yang et al. 2020) (CVPR 20) 19.02 0.577 24.08 0.868
-EnGAN (Jiang et al. 2021) (TIP 21) 17.23 0.543 20.02 0.604
-Uformer (Wang et al. 2022b) (CVPR 22) 18.54 0.577 23.17 0.859
-Restormer (Zamir et al. 2022a) (CVPR 22) 22.27 0.649 25.67 0.827
-MIRNet (Zamir et al. 2022b) (TPAMI 22) 20.84 0.605 24.38 0.864
-SNR-Net (Xu et al. 2022) (CVPR 22) 22.87 0.625 29.44 0.894
-Retinexformer (Cai et al. 2023) (ICCV 23) 24.44 0.680 29.77 0.896
+SID [3] (CVPR 18) 16.97 0.591 23.29 0.703
+RetinexNet [35] (BMVC 18) 16.48 0.578 20.84 0.617
+DeepUPE [28] (CVPR 19) 17.01 0.604 21.70 0.662
+KinD [45] (ACM MM 19) 18.02 0.583 21.95 0.672
+3DLUT [44] (TPAMI 20) 20.11 0.592 21.66 0.655
+DRBN [38] (CVPR 20) 19.02 0.577 24.08 0.868
+EnGAN [12] (TIP 21) 17.23 0.543 20.02 0.604
+Uformer [34] (CVPR 22) 18.54 0.577 23.17 0.859
+Restormer [42] (CVPR 22) 22.27 0.649 25.67 0.827
+MIRNet [43] (TPAMI 22) 20.84 0.605 24.38 0.864
+SNR-Net [37] (CVPR 22) 22.87 0.625 29.44 0.894
+Retinexformer [2] (ICCV 23) 24.44 0.680 29.77 0.896
 Ours-8 25.32 0.802 29.95 0.883
@@
 Table 2: Quantitative comparisons on SID and SDSD datasets. The best result is highlighted in bold, while the second-best result is underlined.
-Dataset and Implementation Details
+4.1 Dataset and Implementation Details
 Datasets.
-We evaluate the performance of our framework across various datasets showcasing noise in low-light image regions, including LOLv1 (Wei et al.
-2018), LOLv2 (Yang et al.
-2021), SID (Chen et al.
-2018), SDSD (Wang et al.
-2021), DICM (Lee, Lee, and Kim 2012), LIME (Guo, Li, and Ling 2016), MEF (Li et al.
-2018), NPE (Wang et al.
-2013), and VV.
-LOLv1 comprises 485 pairs of low/normal-light images for training and 15 pairs for testing, taken under different exposure conditions.
-LOLv2 is divided into LOLv2-real and LOLv2-synthetic subsets.
-LOLv2-real contains 689 pairs for training and 100 for testing, adjusted for exposure time and ISO.
-LOLv2-Synthetic is created by analyzing the illumination distribution of low-light images, comprising 900 pairs for training and 100 for testing.
-SID.
-The subset of the SID dataset captured by a Sony camera is adopted for evaluation.
-There are 2697 short-/long-exposure RAW image pairs.
-The low-/normal-light RGB images are obtained by using the same in-camera signal processing of SID (Chen et al.
-2018) to transfer RAW to sRGB.
-2099 and 598 image pairs are used for training and testing.
-SDSD.
-We adopt the static version of SDSD, captured by a Canon EOS 6D Mark II camera with an ND filter.
-We use 62:6 and 116:10 low-/normal-light video pairs for training and testing on indoor subsets.
-DICM, LIME, MEF, NPE, and VV.
-We evaluate the generalization capability of our models trained on the LOLv1 and LOLv2 datasets by testing them on the DICM, LIME, MEF, NPE, and VV datasets.
-Since these datasets lack ground truth images, evaluation relies on visual comparison and no-reference image quality assessment methods.
+We evaluate the performance of ReDDiT across various datasets showcasing noise in low-light image regions, including LOLv1 [35], LOLv2 [39], SID [3], SDSD [29], DICM [14], LIME [8], MEF [15], NPE [30], and VV.
 Evaluation metrics.
@@
 Methods DICM LIME MEF NPE VV
-LIME (Guo, Li, and Ling 2016) 5.42 3.84 3.73 4.37 3.63
-Zero-DCE (Guo et al. 2020) 3.77 3.61 4.12 6.19 4.81
-EnGAN (Jiang et al. 2021) 3.98 4.49 4.95 3.42 3.93
-KinD++ (Zhang et al. 2021) 6.19 4.24 3.93 3.44 3.72
-SCI (Ma et al. 2022) 3.73 3.52 4.07 6.44 4.05
-DCC-Net (Zhang et al. 2022) 3.84 6.27 6.04 7.97 4.66
-Retinexformer (Cai et al. 2023) 4.47 3.96 4.12 4.28 4.53
-WCDM (Jiang et al. 2023) 6.38 4.62 4.86 4.65 5.39
-GSAD (Hou et al. 2024) 4.35 3.97 4.08 5.35 3.18
+LIME [8] 5.42 3.84 3.73 4.37 3.63
+Zero-DCE [7] 3.77 3.61 4.12 6.19 4.81
+EnGAN [12] 3.98 4.49 4.95 3.42 3.93
+KinD++ [46] 6.19 4.24 3.93 3.44 3.72
+SCI [20] 3.73 3.52 4.07 6.44 4.05
+DCC-Net [47] 3.84 6.27 6.04 7.97 4.66
+Retinexformer [2] 4.47 3.96 4.12 4.28 4.53
+WCDM [11] 6.38 4.62 4.86 4.65 5.39
+GSAD [10] 4.35 3.97 4.08 5.35 3.18
 Ours 3.62 3.45 3.93 3.24 3.00
-Table 3: Quantitative comparisons of different methods on DICM, LIME, MEF, NPE, and VV datasets in terms of NIQE, lower values indicate better performance. The best result is highlighted in bold.
+Table 3: Quantitative comparisons of different methods on DICM, LIME, MEF, NPE, and VV in terms of NIQE. Lower values indicate better performance, with the best result highlighted in bold.
 Figure 5: Quantitative comparisons on the LoLv1 and LoLv2 (Real and Synthetic) datasets with other accelerate methods.
@@
 Our method outperforms other acceleration techniques as well as the original teacher model (in terms of DDIM performance).
-Comparison with State-of-the-Art Methods
+4.2 Comparison with State-of-the-Art Methods
 Results on LOLv1 and LOLv2.
-On LOLv1 and LOLv2, we compare ReDDiT against LIME (Guo, Li, and Ling 2016), RetinexNet (Wei et al.
-2018), KinD (Zhang, Zhang, and Guo 2019), Zero-DCE (Guo et al.
-2020), DRBN (Yang et al.
-2020), KinD++ (Zhang et al.
-2021), EnlightenGAN (Jiang et al.
-2021), MIRNet (Zamir et al.
-2022b), LLFlow (Wang et al.
-2022a), DCC-Net (Zhang et al.
-2022), SNR-Net (Xu et al.
-2022), LLFormer (Wang et al.
-2023b), PyDiff (Zhou, Yang, and Yang 2023), WCDM (Wang et al.
-2023a), and GSAD (Hou et al.
-2024).
-Table 1 presents the quantitative results of various LLIE methods, showcasing our method’s superior performance across all compared metrics, including PSNR, SSIM, and LPIPS.
+On LOLv1 and LOLv2, we compare ReDDiT against LIME [8], RetinexNet [35], KinD [45], Zero-DCE [7], DRBN [38], KinD++ [46], EnlightenGAN [12], MIRNet [43], LLFlow [33], DCC-Net [47], SNR-Net [37], LLFormer [32], PyDiff [48], WCDM [31], and GSAD [10].
+Tab.
+1 presents the quantitative results of various LLIE methods, showcasing our method’s superior performance across all compared metrics, including PSNR, SSIM, and LPIPS.
 New SOTA PSNR results of 28.090, 31.250 and 30.166 can be observed on LoLv1, LoLv2-real and LoLv2-synthetic individually.
@@
 Results on SID and SDSD.
-On SID and SDSD, we compare ReDDiT with SID (Chen et al.
-2018), DeepUPE (Wang et al.
-2019), Uformer (Wang et al.
-2022b), RetinexNet (Wei et al.
-2018), KinD (Zhang, Zhang, and Guo 2019), DRBN (Yang et al.
-2020), EnlightenGAN (Jiang et al.
-2021), MIRNet (Zamir et al.
-2022b), SNR-Net (Xu et al.
-2022), and Retinexformer (Cai et al.
-2023).
-Table 2 presents the quantitative results on the SID and SDSD, indicating its capability of ReDDiT to handle more complex low-light conditions effectively.
+On SID and SDSD, we compare ReDDiT with SID [3], DeepUPE [28], Uformer [34], RetinexNet [35], KinD [45], DRBN [38], EnlightenGAN [12], MIRNet [43], SNR-Net [37], and Retinexformer [2].
+Tab.
+2 presents the quantitative results on the SID and SDSD, indicating its capability of ReDDiT to handle more complex low-light conditions effectively.
 Specifically, our method establishes new SOTA PSNR values of 25.32 dB/29.95 dB on SID and SDSD dataset.
@@
 Results on DICM, LIME, MEF, NPE, and VV.
-As shown in Figure 3, we directly apply the model trained on the LOLv2-synthetic dataset to these unpaired real-world datasets.
-From Table 3, it is evident that ReDDiT outperforms all competitors in terms of NIQE scores, demonstrating its robust generalization ability.
-Clearly, our model adeptly adjusts illumination conditions, effectively enhancing visibility in low-light areas while avoiding over-exposure.
-This successful balance confirms the superior efficacy of our proposed method.
-Comparison with Other Accelerate Methods
+We directly apply the model trained on the LOLv2-synthetic dataset to these unpaired real-world datasets.
+From Tab.
+3, it is evident that ReDDiT outperforms all competitors in terms of NIQE scores, demonstrating its robust generalization ability.
+As shown in Figure 4, our model adeptly adjusts illumination conditions, effectively enhancing visibility in low-light areas while avoiding over-exposure.
+Figure 6: An ablation study conducted to evaluate the choice of $\tilde{x}$.
+The results indicate that reflectance contributes the most to performance during distillation, while low-light images $y$ offer the second-best performance.
+Both reflectance and low-light images $y$ effectively explore a better residual space, leading to significant improvements in performance compared to the original Gaussian space ($x_{0}$ as $\tilde{x}$).
+4.3 Comparison with Other Accelerate Methods
 Figure 7: The ablation study on the main components of our method.
@@
 These findings underscore the significance of ReDDiT’s superior performance, reinforcing its efficacy and the value of employing sophisticated distillation techniques in advancing the SOTA in low-light image enhancement.
-Ablation Study
+4.4 Ablation Study
 The ablation on refinement module.
@@
 In the case of 2-step distillation, Figure 8 illustrates a consistent performance improvement with increasing strength of the refinements.
-On LOLv2-real, the 4-step distillation results even surpass the performance of the 8-step diffusion model and achieve new SOTA results with $w_{refine}=0.8$.
+On LOLv2-real, the 4-step distillation results even surpass the performance of the 8-step diffusion model and achieve new SOTA results with $\omega=0.8$.
 This trend is attributed to the heightened significance of the refinement module in addressing the increasing fitting error of the teacher model as the number of steps decreases.
@@
 This decision ensures that the refinement module effectively mitigates the fitting error of the teacher model, thereby enhancing the performance of the student model during distillation.
+The ablation on the choice of $\tilde{x}$.
+We compare the choice of $\tilde{x}$ with refinement using clean images $x_{0}$, low-light images $y$, the enhanced result from the learnable Retinex method (e.g.
+PairLIE [6]), and our reflectance component, as shown in Figure 6.
+Both reflectance and low-light images $y$ effectively explore a better residual space, leading to significant improvements in performance compared to the original Gaussian space ($x_{0}$ as $\tilde{x}$).
+This suggests that shifting the residual space is crucial for enhancing the performance of the student model in LLIE tasks.
+Furthermore, the PairLIE results, when used for refinement, perform even poorly compared to refinements using clean images $x_{0}$ and low-light images $y$.
+Although its prediction is closer to clean images $x_{0}$, it fails to identify a suitable residual space for the student model to learn from, ultimately failing to address the inference gap.
 The ablation on the auxiliary loss.
@@
 Interestingly, $\mathcal{L}_{per}$ exhibits a more pronounced influence on performance compared to $\mathcal{L}_{pix}$, emphasizing the significance of leveraging supervision signals in the feature space.
-Furthermore, the degradation in performance resulting from ablation of either loss term is minor compared to the substantial degradation observed when the Reflectance-Aware Trajectory Refinement Module is removed.
-This finding highlights the critical role of the Reflectance-Aware Trajectory Refinement Module within our framework.
-Efficiency Comparison
+Furthermore, the degradation in performance resulting from ablation of either loss term is minor compared to the substantial degradation observed when the RATR module is removed.
+This finding highlights the critical role of the RATR module within our framework.
+4.5 Efficiency Comparison
 We evaluate the efficiency of our method in terms of inference time, frames per second (FPS), and number of parameters, comparing it with recent diffusion-based LLIE methods.
-The quantitative results of this comparative analysis are presented in Table 4.
+The quantitative results of this comparative analysis are presented in Tab.
+4.
 Our approach, particularly in its 2-step variant, demonstrates superior performance regarding inference speed, FPS, and parameter efficiency.
@@
 Table 4: The efficiency comparison with other methods. Our method resulted in a fast and lightweight diffusion-based low-light enhancer with excellent performance.
-Limitation and Future work
+4.6 Limitation and Future work
 While ReDDiT has excelled in restoration within 2 steps, single-step restoration is not yet optimal, presenting a notable limitation.
@@
 In our future work, we will continue exploring efficient diffusion-based methods for LLIE.
-Our central focus will be on investigating the potential of a single-step diffusion model and developing a lightweight denoising network.
-Conclusions
-In this paper, we introduce ReDDiT, a significant advancement in efficient diffusion models for low-light image enhancement (LLIE).
-Central to ReDDiT is an innovative distillation technique that integrates deterministic elements and reduces trajectory fitting errors through Reflectance-Aware Trajectory Refinement.
-These advancements enable ReDDiT to preserve the intrinsic structural integrity of images while reducing sampling steps.
-ReDDiT achieves performance comparable to previous methods using just 2 steps and sets new state-of-the-art (SOTA) results in 4 and 8 steps.
+Our focus will be on investigating the potential of a single-step diffusion model and developing a lightweight denoising network.
+5 Conclusions
+In this paper, we theoretically analyze the two main factors contributing to the performance degradation of the diffusion distillation technique.
+Building on this, we introduce ReDDiT, a significant advancement in efficient diffusion models for LLIE.
+Central to ReDDiT is the use of linear exploration in the reflectance-aware residual space, which reduces trajectory fitting errors and the sampling gap.
+These innovations enable ReDDiT to preserve the intrinsic structural integrity of images while minimizing sampling steps.
+ReDDiT achieves performance comparable to previous methods with just 2 steps and sets new SOTA results at 4 and 8 steps.
 Across 10 benchmark datasets, ReDDiT consistently outperforms existing methods.
-This represents a promising step toward real-time diffusion models for LLIE, and our research in this area will continue.
+This marks a promising step toward real-time diffusion models for LLIE, and our research in this area will continue.
+Supplementary Material
+6 Proof of Corollary 1
+Given $\tilde{x}_{0}$, it can be represented into the noisy states according to the forward process of diffusion models as follows:
+$\tilde{x}_{0}=\frac{x_{t}-\sigma_{t}\tilde{\epsilon}}{\alpha_{t}},$ (17)
+Substitute Equation 17 into $\tilde{x}_{s}=\alpha_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}$, we get:
+$\tilde{x}_{s}$ $=\alpha_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}$ (18)
+$=\frac{\alpha_{s}}{\alpha_{t}}x_{t}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\tilde{\epsilon}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u),$
+Given the teacher trajectory $x^{\eta}_{s,u,t}$ defined in the main paper:
+$x^{\eta}_{s,u,t}$ $=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$ (19)
+$+\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}(\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u))$
+$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\epsilon_{\eta}(x_{t},y,t),$
+we can obtain the following equation:
+$\omega x^{\eta}_{s,u,t}+(1-\omega)\tilde{x}_{s}$ (20)
+$=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$
+$+\omega\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}$
+$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\Big{(}\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon}\Big{)}$
+$=\tilde{x}^{\eta}_{s,u,t}.$
+7 Implementation details
+We train our teacher models for 2 million iterations with the framwork of SR3 on the Gaussian flow [24].
+In the distillation stage.
+and 5000 iterations is required.
+We employ the Adam optimizer with a learning rate of $1\times 10^{-4}$, applying linear weight decay.
+To stabilize training, we utilize an exponential moving average (EMA) with a weight of 0.9999 during parameter updates in both the pretrain and distillation stages.
+We adopt a fine-grained diffusion method with $T=512$ steps and implement a linear noise schedule with endpoints set at $1\times 10^{-4}$ and $2\times 10^{-2}$.
+The patch size and batch size are set to 96 and 16, respectively.
+All the experiments are run on a single A100 GPU with 80GB of memory.
+Training our method with a smaller patch size and batch size on a device with less memory is feasible.
+8 Dataset
+LOLv1 comprises 485 pairs of low/normal-light images for training and 15 pairs for testing, taken under different exposure conditions.
+LOLv2 is divided into LOLv2-real and LOLv2-synthetic subsets.
+LOLv2-real contains 689 pairs for training and 100 for testing, adjusted for exposure time and ISO.
+LOLv2-Synthetic is created by analyzing the illumination distribution of low-light images, comprising 900 pairs for training and 100 for testing.
+SID.
+The subset of the SID dataset captured by a Sony camera is adopted for evaluation.
+There are 2697 short-/long-exposure RAW image pairs.
+The low-/normal-light RGB images are obtained by using the same in-camera signal processing of SID [3] to transfer RAW to sRGB.
+2099 and 598 image pairs are used for training and testing.
+SDSD.
+We adopt the static version of SDSD, captured by a Canon EOS 6D Mark II camera with an ND filter.
+We use 62:6 and 116:10 low-/normal-light video pairs for training and testing on indoor subsets.
+DICM, LIME, MEF, NPE, and VV.
+We evaluate the generalization capability of our models trained on the LOLv1 and LOLv2 datasets by testing them on the DICM, LIME, MEF, NPE, and VV datasets.
+Since these datasets lack ground truth images, evaluation relies on visual comparison and no-reference image quality assessment methods.
+9 More experiment results
+Figure 9: Qualitative results on SID and SDSD.
+Patches highlighted in each image by the red and blue boxes indicate that ReDDiT effectively enhances the visibility, preserves the color, reduces noise, and retains finer details in normal light images.
+It can be observed that the visual results of our method align more closely with the ground truth.
+Please zoom in for a clearer view of the image details.
+Figure 10: Qualitative results on DICM, LIME, NPE, MEF, and, VV. Our method effectively enhances the visibility and preserves the color. Please zoom in for a clearer view of the image details.
+9.1 Visual comparison on SID and SDSD dataset
+Due to space constraints in the main manuscript, we only presented quantitative comparisons.
+Here, we offer additional visual comparisons on the SID [3] and SDSD [29] datasets in Figure 9.
+These datasets pose significant challenges for low-light image enhancement due to both low light levels and severe noise degradation.
+Our ReDDiT method demonstrates strong performance in improving brightness and reducing noise on these challenging datasets It can be observed in Figure 9 that our ReDDiT effectively enhances the visibility, preserves the color, improves noise reduction, and retains finer details in normal light images.
+9.2 Visual comparison on Unpair dataset
+Due to space limitations in the main manuscript, we only provided visual comparisons with GSAD [10].
+In Figure 10, we offer additional visual comparisons on DICM, LIME, NPE, MEF, and VV datasets.
+These results demonstrate the generalizability of ReDDiT to real-world low-light image enhancement scenarios and its consistent performance and superiority in preserving image details and improving visibility, making it a reliable solution for real-world applications.